OpenStudy (anonymous):

Please help me medal will be awarded A classroom has a volume of x3 + 6x2 – 1x – 30 ft3 and the height is x – 2 ft. The floor is represented by what polynomial? x2 + 2x – 15 x2 + 8x + 15 x2 + 2x + 8 x2 – 8x + 8

OpenStudy (anonymous):

this is a rectangular paralellopipe volume = area of base * height

OpenStudy (anonymous):

soor hold have to change the question

OpenStudy (anonymous):

by the floor, they mean the area of base \[ x^3+6x^2-x-30=A(x)\times(x-2)\\ A(x)=\frac{x^3+6x^2-x-30}{x-2} \] you can solve this by long-division or by synthetic division